14. Assume that you are told that a variable has a mean of 120 and a variance of 49.

You are asked to construct an interval that includes 75% of the data points.

What would be the upper bound of that interval?

13.

Tammy is a general contractor and has submitted a bid for each of two projects (A and B).

The probability of getting project A is 0.60.

The probability of getting project B is 0.77.

The probability of getting at least one of the projects is 0.89.

What is the probability that she will get both projects?

Question 13 options:

0.67

0.17

0.54

0.60

0.83

0.77

0.48

0.40

0.52

0.55

(hint: You don't know what the distribution "looks like" - it might be bell-shaped or not - you just don't know)

Question 14 options:

198

149

109

134

98

107

none of the above

169
Answer #1

#14.
mean = 120
sd = sqrt(49) = 7

z-value = 1.1503

upper bound = 120 + 1.1503*7 = 128.0521

Using Chebychev's theorem, for k = 2, we get (1 - 1/k^2)% of data within k sigma levels.

Hence upper bound = 120 + 2*7 = 134

#13.
P(A) = 0.6, P(B) = 0.77
P(A or B) = 0.89

P(A and B) = 0.6 + 0.77 - 0.89 = 0.48